OpenStudy (anonymous):

can you find the domain of ln(1-2x)/sqrt(3-x) please!

OpenStudy (anonymous):

Recall that the domain of a function \(f(x)\) is those values of \(x\) for which \(f(x)\) has a value. Hint: Does \(\sqrt{3-x}\) have a value for all values of \(x\)? (We're assuming here that \(x\) is a real number.) What about \(\ln{(1-2x)}\)? Does it have a value for all values of \(x\)?

OpenStudy (anonymous):

you have to put those domains admited by \[\sqrt{3-x} \] and \[\ln (1-2x)\] into a numerical line and find the intersection between them. Always remember taht you cannot divide for 0.